OpenStudy (anonymous):

Need clarification.... 3. Write the first ten terms of the geometric sequence a1 = 3, r = 1.1 to two decimal places. Am I supposed to multiply 3 by 1.1 ten times?

OpenStudy (cwrw238):

a1 = 3 a2 = 3*1.1 a3 = 3 * (1.1)^2 etc

OpenStudy (anonymous):

Thanks! :D

OpenStudy (cwrw238):

yw